OpenStudy (anonymous):

i need help 125/5 / 130/9

OpenStudy (anonymous):

Are you asking\[\frac{ 125 }{ 5 } \frac{ 130 }{ 9 }\]? Is there a + or - or * between them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

/

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

Oh, I see now. Sorry.

OpenStudy (ilovewolf):

125/5 / 130/9 = 1 19/26

OpenStudy (paki):